OpenStudy (anonymous):

Anyone know? The distance d that a spring will stretch varies directly as the force f applied to the spring. If a force of 6 lb is required to stretch the spring 3 in., what force is required to stretch the spring a = 7 in.?

OpenStudy (amistre64):

its a direct variation

OpenStudy (amistre64):

lb = k in k = lb/in

OpenStudy (amistre64):

lb = 2 in

OpenStudy (anonymous):

I don't understand this stuff. Thank You

OpenStudy (amistre64):

well, think of it like this: the force is twice the distance if distance = 3; force = 2*3 = 6 if distance = 7; force = 2*7

OpenStudy (anonymous):

oh ok. That makes sense. I guess I am just very intimidated by word problems.